\begin{thebibliography}{9}
\bibitem{Lucha05:IBSEWEP}W.~Lucha and F.~F.~Sch\"oberl, \emph{J.~Phys.~G:
Nucl.~Part.~Phys.}~\textbf{31}, 1133 (2005) [hep-th/0507281].
\bibitem{RedEq}J.-F.~Laga\"e, \emph{Phys.~Rev.~D} \textbf{45}, 305
(1992); M.~G.~Olsson, S.~Veseli, and K.~Williams,
\emph{Phys.~Rev.~D} \textbf{52}, 5141 (1995) [hep-ph/9503477].
\bibitem{Lucha00/01:IBSE}W.~Lucha, K.~Maung Maung, and F.~F.~Sch\"oberl,
\emph{Phys.~Rev.~D} \textbf{63}, 056002 (2001) [hep-ph/0009185];
``Instantaneous Bethe--Salpeter Equation: (Semi-)Analytical
Solution,'' in \emph{Proceedings of the Int.\ Conf.\ on ``Quark
Confinement and the Hadron Spectrum IV,''} edited by W.~Lucha and
K.~Maung Maung, World Scientific, Singapore, 2002, pp.~340
[hep-ph/0010078]; \emph{Phys.~Rev.~D} \textbf{64}, 036007 (2001)
[hep-ph/0011235]; W.~Lucha and F.~F.~Sch\"oberl,
\emph{Int.~J.~Mod.~Phys.~A} \textbf{17}, 2233 (2002)
[hep-ph/0109165].
\bibitem{Maris}P.~Maris and C.~D.~Roberts, \emph{Phys.~Rev.~C}
\textbf{56}, 3369 (1997) [nucl-th/9708029]; P.~Maris and
P.~C.~Tandy, \emph{Phys.~Rev.~C} \textbf{60}, 055214 (1999)
[nucl-th/9905056]; P.~Maris, ``Continuum QCD and Light
Mesons,''~in \emph{Proceedings of the Int.\ Conference on ``Quark
Confinement and the Hadron Spectrum IV,''} edited by W.~Lucha and
K.~Maung Maung, World Scientific, Singapore, 2002, pp.~163
[nucl-th/0009064].
\bibitem{Lucha05:EQPIBSE}Li Z.-F., W.~Lucha, and F.~F.~Sch\"oberl,
\emph{Mod.~Phys.~Lett.~A} \textbf{21}, 1657 (2006)
[hep-ph/0510372].
\end{thebibliography}

\documentclass[final]{aipproc}\layoutstyle{6x9}\begin{document}
\title{Generalized Instantaneous Bethe--Salpeter Equation and Exact
Quark Propagators}\classification{11.10.St, 03.65.Pm, 03.65.Ge,
12.38.Lg, 12.39.Ki, 14.40.Aq, 14.65.Bt}\keywords{Bethe--Salpeter
equation, relativistic bound state, instantaneous limit, exact
propagator}

\author{Wolfgang Lucha}{address={Institute for High Energy Physics,
Austrian Academy of Sciences,\\Nikolsdorfergasse 18, A-1050
Vienna, Austria\\E-mail: wolfgang.lucha@oeaw.ac.at}}
\author{Franz F.~Sch\"oberl}{address={Department of Theoretical
Physics, University of Vienna,\\Boltzmanngasse 5, A-1090 Vienna,
Austria\\E-mail: franz.schoeberl@univie.ac.at}}\maketitle

Assuming all interactions to be instantaneous \emph{and} the exact
fermion propagator $S$ to be approximately given by ${\rm
i}\,S^{-1}(p)=A(\mathbf{p}^2)\,p\hspace{-1.07ex}/-B(\mathbf{p}^2),$
i.e., with its scalar functions $A$ and $B$ depending only on
$\mathbf{p},$ a three-dimensional reduction of the homogeneous
Bethe--Salpeter equation retaining, in contrast to the Salpeter
equation, the exact propagators (crucial for, e.g., a proper
incorporation of dynamical chiral symmetry breakdown) is
constructed \cite{Lucha05:IBSEWEP}. For spherically symmetric
interactions the resulting bound-state equation reduces to a set
of (coupled) integral equations for the independent radial
components of the bound-state amplitude \cite{RedEq}, which may be
solved by conversion into an equivalent matrix problem
\cite{Lucha00/01:IBSE}. Adopting as light-quark propagator the
solution of the quark Dyson--Schwinger equation found within a
model \cite{Maris} consistent with the QCD axial-vector
Ward--Takahashi identity, a tentative application
\cite{Lucha05:EQPIBSE} to ``pion-like'' pseudoscalar
($J^{PC}=0^{-+}$) mesons considered as quark--antiquark bound
states formed by linear confining interactions of time-component
Lorentz-vector Dirac structure $\gamma^0\otimes\gamma^0$ indicates
that our ``exact-propagator instantaneous formalism'' yields
significantly smaller spacings of its bound-state mass eigenvalues
than those obtained from Salpeter's equation for ``reasonable''
constituent light-quark masses.

\bibliographystyle{aipproc}\end{document}